OpenStudy (anonymous):

(10m9 + 10m6 + 8m4) ÷ 2m2

OpenStudy (anonymous):

answers ; A) 5m9 + 5m6 + 4m4 B) 5m7 + 5m4 + 4m2 C) 5m7 + 10m6 + 8m2 D) 5m7 + 5m4+ 8m4

OpenStudy (anonymous):

B MY MEDAL

OpenStudy (highlyeducated87):

U need an explanation

OpenStudy (anonymous):

divide each term by \(2m^2\) one at a time

OpenStudy (anonymous):

\[\frac{10m^9}{2m^2}=5m^{9-2}=5m^7\] \[\frac{10m^6}{2m^2}=5m^{6-2}=5m^4\] etc

OpenStudy (anonymous):